### v2

# Multivector Reranking in the Era of Strong First-Stage Retrievers

Abstract
[…]
1 Introduction
Transformer-based neural architectures [48] and large pre-trained language models such as BERT [12] have reshaped the field of Information Retrieval (IR), thanks to their ability to encode rich semantic representations of text. These models have enabled a shift from purely lexical matching to representation-based retrieval, where queries and documents are mapped into vector spaces that capture deep contextual meaning. This paradigm has led to a new generation of neural retrievers that substantially improve retrieval effectiveness over traditional term-based models such as BM25 [42]. These models can be broadly categorised into sparse and dense encoders. Sparse encoders, such as Splade [21], produce high-dimensional representations, where only a small fraction of components are non-zero. In contrast, dense encoders represent each query or document as a dense vector in a low-dimensional space that captures global semantic similarity. They can be further divided into single-vector and multivector architectures. Single-vector encoders [25, 49, 50] map an entire passage to a single dense vector, providing compact document-level semantics but limiting fine-grained query–document interactions. Differently, multivector models such as ColBERT [26] encode each query and document as a set of token-level vectors. In this framework, similarity between a query and a document is computed via late interaction through the max similarity operator, $MaxSim(q,D)=\sum_{i=1}^{n_{\mathtt{q}}}\max_{j=1,\dots,n_{\mathtt{D}}}\langle q_{i},D_{j}\rangle$, which lets every query token interact with every document token [26]. These models achieve state-of-the-art effectiveness because they preserve fine-grained token interactions. However, performing retrieval over a full multivector collection is computationally infeasible. For this reason, all scalable multivector systems follow a gather-and-refine strategy [43, 38, 2, 45]: 1) a gather phase that selects a relatively small set of candidate documents; 2) a refine phase that applies full $MaxSim$ scoring on these candidates to produce the final ranking.
Most existing systems perform the gather step at the token-embedding level. As already noted in [29], this design has two fundamental drawbacks: 1) the index cardinality (the number of vectors to search) grows with the number of tokens, which is typically one or two orders of magnitude larger than the number of documents, leading to significantly higher search-space complexity and slower queries, and 2) token-level similarity alone does not predict document-level relevance well: $MaxSim$ depends on the joint interaction of all query tokens with all document tokens, and the final score depends on many different document tokens. The token-based gather step is therefore both inefficient, because it must retrieve and evaluate many more candidates, and often ineffective, because it fails to identify the documents that would score highest under full $MaxSim$. To mitigate this, XTR [29] redefines the training objective to make token retrieval itself more predictive of document-level relevance, partially closing the training–inference gap.
Through a reproducibility study (Section 4.1), we observe that state-of-the-art multivector methods exhibit limited efficiency. Despite their effort to optimize the gather phase, this step remains fundamentally inadequate. As emerges in recent works [38, 45, 2], token-level gathering not only incurs high computational costs but also lacks the selectivity required for a compact candidate set, often returning hundreds or thousands of documents that must undergo expensive full MaxSim scoring. Driven by this observation, in contrast to the approach taken by XTR [29], we take a different direction: we propose to replace the token-level gather step entirely with a document-oriented gather strategy, where candidates are selected based on single-vector document representations. This design naturally aligns with the classical two-stage retrieval architecture. In literature, BM25 and similar lexical bag-of-words models have been widely used for first-stage retrieval. However, the limited semantic coverage of BM25 comes at a price: if the initial candidate set omits a large fraction of relevant documents, no reranking strategy can fully recover them. Indeed, a recent work shows that the effectiveness achieved by reranking BM25 candidates using multivectors is limited [33].
We claim that recent advances in first-stage retrievers based on learned sparse representations make this shift particularly effective, yielding more semantically coherent candidate sets while drastically reducing retrieval cost. To support this, Figure 1 (left) compares the recall at various cut-offs $\kappa$ between BM25 and a learned sparse representation obtained with the Splade CoCondenser EnsembleDistill model on the Ms Marco-v1 passage dataset. The recall gap is substantial. To compensate for the recall deficit of BM25, pipelines typically increase the cut-off $\kappa$ to hundreds or even thousands of documents, which in turn inflates reranking cost. From Figure 1 (right), it is clear that reranking time grows rapidly with the number of candidates, making larger cut-offs increasingly inefficient.
[…]
Even if with large $\kappa$ BM25 manages to reach quite high recall, Figure 1 (center) shows that the resulting MRR@10 after reranking is still much lower. This indicates that BM25 misses many semantically relevant candidates and that increasing $\kappa$ introduces noisy documents whose scores hurt reranking quality.
In this paper, we analyse a modern two-stage pipeline where both stages use neural representations: a learned sparse retriever for the first stage and a dense multivector late-interaction reranker. Our contributions are:
  • We reproduce several state-of-the-art multivector retrieval methods on two publicly available datasets, providing a clear picture of the current multivector retrieval field.
  • We introduce a hybrid architecture that couples LSR with multivector reranking. Experiments on two public datasets show that our hybrid pipeline consistently achieves one order of magnitude speedup at comparable or superior effectiveness, reaching up to $24\times$ over the best-performing competitor under similar memory and encoder settings.
  • To enable full reproducibility of our findings, we provide a public—end-to-end—implementation of our proposed retrieval pipeline. Our implementation is available on GitHub.
2 Related Work
Multivector Retrieval via Late-Interaction. Late-interaction models such as ColBERT [26] and Xtr [29] represent each document as a set of token-level vectors and compute relevance through the $MaxSim$ operator, which allows every query token to interact with every document token. This fine-grained interaction delivers state-of-the-art effectiveness but makes exhaustive retrieval computationally infeasible.
Efficient Multivector Retrieval Systems. To address this inefficiency, early systems implemented token-level gather, which retrieves documents based on token embedding scores. Several works have proposed mechanisms to make multivector retrieval efficient by improving either the gather or the refine step. DESSERT [15] improves the gather phase using hashing-based approximate scoring, which leads to a more efficient refine phase. Plaid [43] clustered token embeddings into centroids, enabling inverted-file filtering before refinement, while ColBERTv2 [44] introduced residual compression and improved training. CITADEL [36] proposed conditional token interaction via dynamic lexical routing, allowing each query to estimate document relevance by focusing on a subset of relevant document tokens. Emvb [38] further optimized ColBERT-style retrieval through bit-vector prefiltering, SIMD acceleration, and PQ-based [24] refinement, achieving up to $2.8\times$ faster retrieval than Plaid, with minimal quality loss. Warp [45] explored a complementary path, combining ColBERTv2 ’s centroid interaction with XTR’s lightweight architecture [29], achieving $3\times$ speed-ups over ColBERTv2/Plaid. Igp [2] introduced an incremental greedy probing strategy over proximity-graph indexes, which incrementally expands search regions and finds high-quality candidates with fewer probes. A different line of work, MUVERA [13], reformulates multivector retrieval as single-vector retrieval by partitioning the token space and aggregating embeddings into compositional vector codes, achieving high efficiency but at the cost of flexibility and fine-grained scoring.
Learned Sparse Retrieval. Learned Sparse Retrieval (LSR) approaches bridge the gap between lexical and dense models by generating sparse term-weight representations that capture contextual semantics beyond exact token matching [1, 9, 10]. Early works such as UniCOIL-T5 [31, 32] expanded document text with contextual terms predicted by DocT5Query [41], enabling lexical-style indexing with neural enrichment. A cornerstone in this family is Splade [21], which introduced sparsity-inducing regularization and log-saturation on term weights to produce highly sparse, interpretable representations while preserving semantic coverage. Subsequent variants refined its efficiency and regularization objectives [18, 28, 20], further improving effectiveness and scalability. More recently, sparse retrieval has been extended beyond text to multimodal settings [39]. A promising line of investigation for LSR as first-stage retrieval is the “Inference-Less” paradigm, that removes the query encoder and computes query weights from fixed, precomputed lookup tables. Early implementations adopted the simplest strategy of assigning a uniform score of 1 to all query terms [18, 28]. Subsequent work introduced more sophisticated scoring mechanisms like integrating an Inverted Document Frequency (IDF) term into the loss function and use term IDF as query scores [23]. Shen et al. introduced a $l_{0}$ regularization to balance sparsity and avoid documents to collapse to all zero components [46]. Learned Inference-less Sparse Retrieval (Li-Lsr[37] learns an optimal lookup table $\text{term}\rightarrow\text{score}$ at training time by projecting the static word embeddings into a scalar value by means of a linear layer.
Hybrid Architectures. Other research explicitly combines sparse and dense models. For instance, BGE-M3 [8] encoder offers support to dense, sparse and multivector retrieval seamlessly, while SparseEmbed [27] extends Splade with dense embeddings. Approaches like SLIM [30] and Splate put together sparse and multivector representations. Splate [17] learns a sparse adapter on top of frozen ColBERT embeddings via distillation from Splade, enabling ColBERT-style reranking with sparse, interpretable candidate generation.
Efficient Engines for Large-scale Retrieval. Modern retrieval pipelines rely on efficient Approximate Nearest Neighbors (ANN) search or inverted-index engines to support fast retrieval. kANNolo [11] is a research-oriented ANN library that supports both dense and sparse retrieval. Its sparse-HNSW implementation has shown [11] to outperform the winners of the “Sparse Track” at NeurIPS 2023 Big ANN challenge [47], thus representing the state-of-the-art among graph indexes for sparse retrieval. Complementary to graph-based ANN approaches, another family of systems leverages inverted indexes optimized for LSR, including Block-Max Pruning [34], Dynamic Superblock Pruning [7], and Seismic [4, 5, 6]. Seismic represents the current state-of-the-art in LSR. It partitions each inverted list into geometrically cohesive blocks represented by summary vectors. At query time, it ranks blocks via their summaries and fully evaluates only those above a relevance threshold, greatly reducing computation while maintaining accuracy.
3 Methodology
In this work, we aim to reproduce and analyze the main multivector retrieval methods proposed in the recent literature. Our goal is to clarify their relative strengths, weaknesses, and trade-offs between efficiency and effectiveness under a unified evaluation framework. Moreover, we extend experimental evaluation with our hybrid two-stage retrieval architecture that integrates learned sparse retrieval for candidate generation and multivector reranking for refinement.
3.1 Two-stage Retrieval
In our two-stage framework, we employ Learned Sparse Retrieval (LSR) methods as the first-stage retrievers to collect $\kappa$ candidate documents. The choice of LSR over dense single-vector representations has many motivations: 1) LSR generally offers similar performance to dense retrieval with lower memory overhead; 2) LSR can leverage inferenceless methods to remove the need for a query encoder; 3) LSR supports efficient inverted index retrieval [4, 34]. The candidate set is then re-ranked using a multivector representation, from which we extract the top $\kappa_{f}=10$ results.
Encoders. We adopt the Splade CoCondenser model as our sparse encoder [19] (MRR@10 on MsMarco $38.3$, Success@5 on LoTTE $69.0$), and ColBERTv2 as our dense multivector encoder [44]. Both methods are reference baselines for sparse and multivector encoders. For the inference-less sparse embeddings, we adopt the Big version of the Li-Lsr model proposed in [37] (MRR@10 on MsMarco $38.8$, Success@5 on LoTTE $65.7$).
First-Stage Retrievers. kANNolo and Seismic are the state-of-the-art solutions for sparse graphs and inverted index for retrieving over LSR representations [11, 4, 35, 3]. Hence, we pick them as our retrieval engines. In the inference-less setting, we rely exclusively on Seismic, since, as noted in [4], graph-based approaches tend to be less effective when the query is much sparser than the documents.
[…]
When employing Seismic as the first-stage retriever, we emulate the same pipeline by running a grid-search over Seismic’s parameters to identify the fastest configurations achieving given accuracy thresholds. For each of these configurations, we rerank progressively larger candidate sets to assess the efficiency–effectiveness trade-off. We use our newly introduced RerankIndex to rerank the results of Seismic search.
Multivector Compression. In our two-stage scenario, we evaluate four different compression configurations for multivector embeddings, representing a broad spectrum of memory–efficiency–effectiveness trade-offs:
Half-precision: This setting provides maximum accuracy as it uses 16-bit floating-point per vector dimension. As ColBERT embeddings have $128$ dimensions, which results in 256 bytes per token embedding, this representation suffers a high memory consumption.
OPQ: Optimized Product Quantization [22] is a variant of PQ [24] that rotates the subspaces to improve compression. We use $M=64$ subspaces, resulting in $64$ bytes per token embedding. This setting reduces memory by $4\times$ but increases reranking latency due to subspace-wise distance precomputation.
MOPQ: Multivector-OPQ uses i) $\kappa$-Means to compute first-level centroids, and ii) OPQ with $M=32$ to compress the residuals between the original vector and the centroids. This approach takes $4$ bytes to store the centroid id, and 32 bytes for OPQ-encoded residuals, resulting in 36 bytes per token (a $7\times$ reduction).
JMPQ: Jointly-optimized Multivector Product Quantization [16], is a supervised version of Multivector-PQ that jointly optimizes both levels and the query encoder to minimize the reconstruction error. The model we used was trained on MsMarco. We test two configurations: 16 and 32 subspaces, resulting in 20 and 36 bytes per token ($12.8\times$ and $7\times$ reductions, respectively).
The compressed configurations require an additional preprocessing phase. OPQ involves relatively fast training since it operates on small subspaces, whereas both MOPQ and JMPQ employ $\kappa$-means clustering at the first level, introducing a notable preprocessing overhead. JMPQ further adds complexity, as it requires end-to-end joint optimisation during training.
Reranking Optimizations. We propose two novel lightweight optimization strategies to improve reranking efficiency. Candidates Pruning (CP): we truncate the candidate list using a threshold $\alpha$ based on the difference between the $\kappa_{f}$-th candidate score and subsequent scores. A sharp drop in first-stage scores suggests a general relevance decline, allowing us to safely discard lower-ranked documents. More precisely, let $t$ be the first-stage score of the $\kappa_{f}$th candidate; if, for the first-stage score $s$ of a candidate it holds $s<(1-\alpha)t$, then we discard the current candidate and all the other candidates below it. Early Exit (EE): during reranking, if the top-$\kappa_{f}$ results remain unchanged for $\beta$ consecutive candidates, we stop reranking early, speculating on the fact that successive candidates are unlikely to improve the final ranking. In particular, if for $\beta$ steps the set of $\kappa_{f}$ best results so far is not updated after computing late interaction score, then stop reranking and return the current set of best results.
Parameters Selection. When building HNSW with kANNolo, we experimentally verify that number of neighbors $M=32$ and construction precision $ef_{c}\in\{1500,2000\}$ yield optimal results respectively on LoTTE and MsMarco. At search time, the search expansion factor ($ef_{s}$) tunes the efficiency-effectiveness tradeoff. Thus, we perform a complete grid search over the set $\{ef_{s},\kappa,\alpha,\beta\}$.
For Seismic, we first tune its construction and search parameters to reach Accuracy@50 (the fraction of true nearest neighbors contained within the top-50 approximate retrieval results) levels in $[0.90,0.99]$ with $0.01$ steps. For each configuration, we then perform offline reranking at different cut-offs of $\kappa\leq 50$ and sum the first- and second-stage latencies.
The reranking hyperparameters tested are: $\kappa\in\{15,20,25,30,35,40,50\}$, $\alpha\in\{0.015,0.025,0.050,\text{{off}}\}$, $\beta\in\{2,3,4,\text{{off}}\}$, where “off ” means that the optimization (CP or EE) is disabled.
3.2 Competitors
We report the detailed experimental configuration of the evaluated methods. Whenever possible, we rely on the pre-built indexes releases by the authors, to ensure a fair reproduction of their results and avoid crafting ourselves sub-optimal configurations.
EMVB. Indexes are publicly available on the official GitHub repository. [^1] https://github.com/CosimoRulli/emvb While the authors provide the hyperparameters to replicate the original results, we perform a grid search to validate this approach further. The specific ranges for each parameter can be found in our code. EMVB is written in C++.
IGP. Indexes are also publicly available on their GitHub repo. [^2] https://github.com/DBGroup-SUSTech/multi-vector-retrieval The authors in the paper indicate specific ranges for both search parameters. In particular, they recommend $\phi_{pb}\in\{1,2,4,8,16,32,64\}$ and $\phi_{ref}\in\{1,2,5,10,20,30,40,50\}\cdot k_{f}$. IGP is written in C++.
WARP. Warp’s authors did not test their method on MsMarco and they employ LoTTE only on XTR’s encodings, so using Warp’s code we created our own encodings and built our own indexes. For the construction parameters, the authors suggest $b\in\{2,4\}$ bits per component to encode the residual embeddings. We choose $b=2$ in order for us to have a direct comparison against Igp (2 bits), Emvb (MOPQ32, resulting in 2 bits per component) and our MOPQ32/JMPQ32 settings. Other construction parameters, like the number of $\kappa$-means centroids, are empirically computed based on the characteristics of the dataset. Two parameters control the search quality in Warp: the number of clusters to decompress per query token $n_{probe}$ and the threshold on the cluster size used for $\text{WARP}_{\text{SELECT}}$ $t^{\prime}$. From their experiment, they find that Warp is robust to variations of $t^{\prime}$ and they choose to compute it empirically basing on dataset’s size. Thus, for our grid search, we let $n_{probe}$ vary in the range $\{1,2,4,8,16,32,64\}$ as they suggest.
4 Experiments
We conduct experiments on the MsMarco passages dataset [40] for the in-domain evaluation and on LoTTE-pooled [44] for the out-of-domain evaluation. The MsMarco dataset is a collection of $8.8$ M passages; we use the 6,9806{,}980 queries in the dev.small to assess the effectiveness of our approaches, measured in terms of MRR@10. LoTTE-pooled contains approximately $2.4$ M passages drawn from multiple domains (e.g., search, forum, Wikipedia, news, and StackExchange). We use the search/dev split comprising 2,9312{,}931 queries. We generate the multivector embeddings for both collections with the original ColBERT code, using the ColBERTv2 checkpoint. [^3] https://github.com/stanford-futuredata/ColBERT Optimized Product Quantization is trained using the Faiss library [14], while for JMPQ we use the original code [16].
Hardware Details. Our experiments were executed on a server equipped with 512 GiB of RAM and 2 Intel Xeon Silver 4314 clocked at 22.44GHz, with $64$ total threads between physical and hyper-threaded ones. Index construction is performed using multi-threading to exploit available CPU resources fully. In contrast, all search operations are executed sequentially on a single core to ensure fair and consistent performance comparisons across different indexing methods and settings. Our methods are implemented in Rust (cargo 1.92.0-nightly) and compiled in release mode to enable CPU-specific optimizations.
4.1 Reproducibility and Comparison with the State of the Art
Tables 1 and 2 compare our reranking methods against Emvb, Warp, and Igp in terms of memory usage (bytes per token vector), effectiveness (MRR@10 or Success@5), and retrieval time (ms) on Ms Marco-v1 and LoTTE-pooled, respectively.
For our first stage, we consider three variants: i) Double-encoder - kANNolo, where we employ the end-to-end RerankIndex implemented in kANNolo and the first stage relies on Splade CoCondenser representations; ii) Double-encoder - Seismic, where the first stage still relies on Splade CoCondenser as a sparse encoder, but we use Seismic for first-stage retrieval; iii) Li-Lsr - Seismic, that replaces Splade CoCondenser with the Li-Lsr Big model. The reranking is tested with the half-precision baseline, JMPQ16/32 (on MsMarco) and MOP32 (on LoTTE).
Table 1: Retrieval time (ms) on Ms Marco-v1 (- indicates not reached). Bytes per vector for Warp are estimated from index size.
Bytes MRR@10
0.390 0.392 0.394 0.396 0.398 0.399
Warp $35$ $98.7$ $98.7$ $136.3$ - - -
Igp $36$ $72.8$ $110.0$ - - - -
Emvb (JMPQ16) 2020 $56.0$ $56.7$ $61.4$ - - -
Emvb (JMPQ32) $36$ $56.0$ $56.0$ $56.4$ $57.3$ $59.4$ $67.1$
Double-encoder - kANNolo
   w/ Half-precision $256$ $1.3$ $1.3$ $1.4$ $1.7$ $1.8$ $1.9$
   w/ JMPQ16 2020 $2.4$ $2.9$ $4.2$ $6.6$ - -
   w/ JMPQ32 $36$ $2.8$ $3.0$ $3.4$ $4.2$ $5.5$ $8.1$
Double-encoder - Seismic
   w/ Half-precision $256$ $0.5$ $0.5$ $0.5$ $0.7$ $1.5$ -
   w/ JMPQ16 2020 $1.0$ $1.0$ $1.3$ - - -
   w/ JMPQ32 $36$ $1.5$ $1.5$ $1.5$ $1.5$ $2.4$ $2.5$
Li-Lsr - Seismic
   w/ Half-precision $256$ $1.5$ $1.5$ $1.7$ $2.3$ $5.4$ -
   w/ JMPQ16 2020 $2.3$ $2.9$ $5.8$ - - -
   w/ JMPQ32 $36$ $2.4$ $2.4$ $2.4$ $2.6$ $5.7$ -
Table 2: Retrieval time (ms) on LoTTE-pooled (- indicates not reached). Bytes per vector for Warp are estimated from index size.
Bytes Success@5
$67.0$ $67.5$ $68.0$ $68.5$ $69.0$ $69.5$
Warp $35$ $49.9$ $49.9$ - - - -
Igp $36$ $39.0$ $48.7$ - - - -
Emvb (MOPQ32) $36$ $54.0$ $54.7$ $55.4$ $59.2$ $74.2$ -
Double-encoder - kANNolo
   w/ Half-precision $256$ $1.7$ $1.8$ $2.1$ $2.4$ $3.0$ $4.0$
   w/ MOPQ32 $36$ $4.0$ $4.0$ $4.3$ $4.7$ $5.8$ $10.1$
Double-encoder - Seismic
   w/ Half-precision $256$ $1.3$ $1.3$ $1.3$ $1.3$ $1.3$ $1.4$
   w/ MOPQ32 $36$ $3.1$ $3.1$ $3.1$ $3.1$ $3.1$ -
Li-Lsr - Seismic
   w/ Half-precision $256$ $3.3$ $3.3$ $3.3$ $3.3$ $3.7$ -
   w/ MOPQ32 $36$ $5.0$ $5.0$ $5.0$ $5.4$ - -
In both tables, we report the cost of storing a term vector for all the multi-vector methods (Emvb, Igp, Warp). For our reranking methods, observe that storing the sparse embeddings induces additional memory overhead. On MsMarco, the Splade CoCondenser embeddings have a size of approximately 4 GB, whereas the inferenceless variant requires about 13 GB. On LoTTE, the corresponding sizes are approximately 1.5 GB and 4.2 GB for Splade CoCondenser and inferenceless data, respectively. For completeness, the size of the multivector datasets is 22.5 GB and 13.5 GB, respectively for JMPQ32 and JMPQ16 on MsMarco, and 10 GB for MOPQ32 on LoTTE. The storage required by the index structures themselves is negligible and therefore omitted for both sparse and multivector indexes.
Comparative Results Analysis. On MsMarco (Table 1), Emvb offers superior performance than both Igp and Warp, with both lower latencies and higher peak effectiveness. On LoTTE (Table 2), Emvb has slightly higher latencies than Warp and Igp but still yields higher peak effectiveness. Overall, our reproduced performance aligns well with the results reported in the original Emvb paper. Warp, though originally designed for Xtr, performs comparably to Igp, confirming the generalization capability to ColBERTv2 that was noted by its authors in the appendix of [45].
On MsMarco, Igp demonstrates effectiveness levels consistent with those reported in the original paper [2], although its reported latencies are significantly lower than those reported for Emvb; in our experiments, however, we observe the opposite trend, with Emvb consistently outperforming Igp. The absolute differences between our measured times and those reported in the original paper may be due to different hardware configurations, while the reversed trend likely arises from the authors not using the optimized indexes provided in Emvb’s official repository nor its original index-building code. [^4] The authors implemented their own index building https://tinyurl.com/3yd8fjam. We highlight that in the Igp original paper, the authors evaluated Igp on different metrics, specifically on MRR and Recall, while we attain to use the LoTTE official metric [44], i.e., Success@5.
Across both datasets, all three of our configurations are approximately one order of magnitude faster than the competitor methods at comparable memory budgets. This advantage is due to the small set of candidates we need to rerank ($\leq 50$ in all our settings) and the efficient first-stage retrievers we employed. Because our reranking stage operates on a small candidate set—and our optimizations (CP, EE) further reduce the number of full evaluations—the pipeline’s bottleneck shifts to the first-stage retrieval. Consequently, in the double-encoder setting, Seismic provides significant benefits over kANNolo, especially at higher effectiveness levels where latency becomes dominated by the first stage. The inferenceless configuration exhibits higher latency due to heavier vectors and the larger number of non-zero components, which increase index size and slow similarity computations. However, this setup is directly comparable to the competitors since it uses only one encoder at query time. Under this fair comparison (i.e., with only one encoder involved), the inferenceless compressed setting achieves speed-ups of $10\times$$24\times$ on MsMarco and $7.8\times$$11\times$ on LoTTE relative to the fastest competing method.
Reproducibility Remarks. Overall, our reproduced results are often consistent with the original findings. Emvb, though preceding Igp and Warp, shows stable and competitive performance in both latency and effectiveness, confirming the authors’ reported results. Warp generalises beyond its native Xtr setup. However, a direct comparison to Emvb was missing in prior work, and we close this gap with our evaluation. Igp reproduces the reported effectiveness on MsMarco but exhibits higher latencies relative to Emvb, likely due to the use of different indexes and modified build procedures. Moreover, its evaluation does not include LoTTE’s official metric (Success@5).
Table 3: Evaluation of reranking optimizations. End-to-end (1st + 2nd stage) retrieval time (ms) on MsMarco.
MRR@10
0.400 0.402 0.404 0.405
Half-precision $2.5$ $3.7$ $6.7$ -
   + CP $2.2$ $3.2$ $5.7$ $9.8$
   + EE $2.2$ $3.2$ $5.6$ -
   + CP + EE $2.1$ $2.8$ $5.6$ -
4.2 Impact of Quantization and Reranking Optimizations
We perform an ablation study on the MsMarco dataset to analyze the effect of multivector quantization and our proposed reranking optimizations, i.e. Candidate Pruning (CP) and Early Exit (EE). Table 3 and Figure 2 report MRR@10 versus reranking latency for all tested configurations on the MsMarco dataset, where on first stage we employ Splade CoCondenser representations.
Table 3 isolates the Half-precision configuration. This setup defines the upper bound in retrieval quality, reaching MRR@10 above $0.40$, but at a very high memory cost ($256$ Bytes per token embedding). Latency ranges between 2.5 and 9.8 ms per query, depending on the MRR target. Both CP and EE reduce latency by 10–20% with no impact on quality, while their combination (CP+EE) offers the best balance, achieving up to $2.1$ ms per query at similar MRR.
The plots in Figure 2 compare all five compression schemes, i.e., Half-precision, OPQ64, MOPQ32, JMPQ32 and JMPQ16, under different optimization settings. All compressed variants exhibit slightly lower MRR (up to $\approx$0.399) but deliver substantial memory savings, from $4\times$ (OPQ64) to $12.8\times$ (JMPQ16) relative to the Half-precision baseline.
Figure 2: Evaluation of quantization schemes and optimizations. End-to-end (1st + 2nd stage) retrieval time (ms) on MsMarco.
OPQ64 maintains strong accuracy but suffers from the highest latency (up to 13.5 ms) due to the computation of 64 distance tables per query token. MOPQ32 achieves similar performance at roughly half the memory footprint (36 Bytes per embedding), although its peak effectiveness is slightly lower. JMPQ32 and JMPQ16 deliver the best overall trade-offs: they retain most of the effectiveness while improving efficiency and compressing memory by $7\times$$12.8\times$. Across all quantization settings, CP and EE consistently improve efficiency with no losses in MRR. The best results, as in the Table 3, are obtained combining CP and EE, achieving a maximum speed-up of $1.8\times$.
Overall, these results confirm that: (i) quantization can substantially reduce memory with limited effectiveness loss, (ii) CP and EE provide substantial efficiency gains at no cost in effectiveness, and (iii) JMPQ achieves the best overall efficiency–effectiveness balance, but it needs supervised training; MOPQ and OPQ still offer valid alternatives.
5 Conclusion
We revisited recent multivector retrieval methods within a unified, fully reproducible framework, reproducing and comparing three state-of-the-art systems—Emvb, Warp, and Igp—under identical hardware and evaluation settings. Our reproduction confirms that, despite substantial engineering optimizations, these token-level Gather & Refine pipelines exhibit high query latency in practice.
Although we do not decompose the internal costs of the gather and refine phases, this behavior is consistent with the design of token-level retrieval, which relies on expensive token-level indexes and retrieves hundreds or thousands of candidates for full-scoring, as documented in prior work.
Motivated by this observation, we explored an alternative design in which the gather phase operates at the document level using learned sparse retrieval, and multivector scoring is applied only as a reranking step. Our benchmarks show that this two-stage pipeline requires only 20205050 candidates for reranking to match or exceed the effectiveness of gather-based multivector systems, while substantially reducing query latency. Under comparable memory and encoder settings, this approach achieves speed-ups of up to $24\times$ over the fastest reproduced baselines. Our experimental analysis further investigates quantization and novel pruning and early-exit strategies, showing that quantization effectively reduces memory usage and that adaptive reranking consistently improve efficiency.
6 Acknoledgments
This work was partially supported by the European Union’s Horizon 2020 Research and Innovation Staff Exchange programme under the Marie Skłodowska-Curie grant agreement No. 872539, by the Horizon Europe RIA “Extreme Food Risk Analytics” (EFRA), grant agreement No. 101093026, by the PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - “FAIR - Future Artificial Intelligence Research” - Spoke 1 “Human-centered AI” funded by the European Commission under the NextGeneration EU program, and by the MUR-PRIN 2022 “Algorithmic Problems and Machine Learning”, grant agreement n. 20229BCXNW.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v3

@@
 Double-encoder - kANNolo
-   w/ Half-precision $256$ $1.3$ $1.3$ $1.4$ $1.7$ $1.8$ $1.9$
+   w/ Half-precision $256$ $2.3$ $2.8$ $3.0$ $4.1$ $6.2$ $8.7$
    w/ JMPQ16 2020 $2.4$ $2.9$ $4.2$ $6.6$ - -
@@
 Moreover, its evaluation does not include LoTTE’s official metric (Success@5).
-Table 3: Evaluation of reranking optimizations. End-to-end (1st + 2nd stage) retrieval time (ms) on MsMarco.
-MRR@10
-0.400 0.402 0.404 0.405
-Half-precision $2.5$ $3.7$ $6.7$ -
-   + CP $2.2$ $3.2$ $5.7$ $9.8$
-   + EE $2.2$ $3.2$ $5.6$ -
-   + CP + EE $2.1$ $2.8$ $5.6$ -
 4.2 Impact of Quantization and Reranking Optimizations
@@
 Candidate Pruning (CP) and Early Exit (EE).
-Table 3 and Figure 2 report MRR@10 versus reranking latency for all tested configurations on the MsMarco dataset, where on first stage we employ Splade CoCondenser representations.
-Table 3 isolates the Half-precision configuration.
-This setup defines the upper bound in retrieval quality, reaching MRR@10 above $0.40$, but at a very high memory cost ($256$ Bytes per token embedding).
-Latency ranges between 2.5 and 9.8 ms per query, depending on the MRR target.
-Both CP and EE reduce latency by 10–20% with no impact on quality, while their combination (CP+EE) offers the best balance, achieving up to $2.1$ ms per query at similar MRR.
-The plots in Figure 2 compare all five compression schemes, i.e., Half-precision, OPQ64, MOPQ32, JMPQ32 and JMPQ16, under different optimization settings.
-All compressed variants exhibit slightly lower MRR (up to $\approx$0.399) but deliver substantial memory savings, from $4\times$ (OPQ64) to $12.8\times$ (JMPQ16) relative to the Half-precision baseline.
+Figure 2 reports MRR@10 versus search latency for all tested configurations on the MsMarco dataset, where on first stage we employ Splade CoCondenser representations.
+The figure compare all five compression schemes, i.e., Half-precision, OPQ64, MOPQ32, JMPQ32 and JMPQ16, under different optimization settings.
+JMPQ32 shows no loss in quality despite a $7\times$ memory reduction compared to the Half-precision setting.
+The other compressed variants exhibit slightly lower MRR but deliver substantial memory savings, ranging from $4\times$ (OPQ64) to $12.8\times$ (JMPQ16) relative to the baseline.
 Figure 2: Evaluation of quantization schemes and optimizations. End-to-end (1st + 2nd stage) retrieval time (ms) on MsMarco.
@@
 JMPQ32 and JMPQ16 deliver the best overall trade-offs: they retain most of the effectiveness while improving efficiency and compressing memory by $7\times$$12.8\times$.
-Across all quantization settings, CP and EE consistently improve efficiency with no losses in MRR.
-The best results, as in the Table 3, are obtained combining CP and EE, achieving a maximum speed-up of $1.8\times$.
+Regarding the reranking optimizations, CP and EE consistently improve efficiency across all quantization settings, achieving a maximum speed-up of $1.8\times$.
+CP is the most reliable optimization; by relying on uncompressed first-stage scores, it consistently reaches the same MRR as its unoptimized counterpart.
+Conversely, EE is occasionally too aggressive, preventing the algorithm from reaching top accuracy; this also affects the combined EE + CP setting.
 Overall, these results confirm that: (i) quantization can substantially reduce memory with limited effectiveness loss, (ii) CP and EE provide substantial efficiency gains at no cost in effectiveness, and (iii) JMPQ achieves the best overall efficiency–effectiveness balance, but it needs supervised training; MOPQ and OPQ still offer valid alternatives.